OpenStudy (rowanmartinfan):

@jcpd910 >>> The equation T^2 = A^3 shows the relationship between a planet’s orbital period, T, and the planet’s mean distance from the sun, A, in astronomical units, AU. If planet Y is k times the mean distance from the sun as planet X, by what factor is the orbital period increased? <<<

OpenStudy (anonymous):

I'm sorry, I have a massive problem picturing things like this in my head, so I need an image. :/

OpenStudy (anonymous):

If \[T^{2}=A^{2}\] lets assume that planet X has A astronomical units for a mean distance from the sun. That means that planet Y has a mean distance of KA. So\[T^{2}=(KA)^{2}=K^{2}A^{2}\]for planet Y.

OpenStudy (anonymous):

Oops I thought it was T^2, the only thing that changes is\[T^{2}=(KA)^{3}=K^{3}A^{3}\]

OpenStudy (rowanmartinfan):

T^2 = a^3

OpenStudy (anonymous):

Yes I just typed it wrong...again.

OpenStudy (rowanmartinfan):

so the ans is

OpenStudy (anonymous):

What would the orbital period be for planet X in this situation?

OpenStudy (rowanmartinfan):

hm 1/2

OpenStudy (anonymous):

You would solve for T in the equation \[T^{2}=A^{3}\]

OpenStudy (rowanmartinfan):

ooh 2/3

OpenStudy (anonymous):

If \[T^{2}=A^{3}\]then to solve for T, we would take the square root of both sides.

OpenStudy (rowanmartinfan):

i still got 2/3

OpenStudy (anonymous):

Show me your work. Your answer should be in terms of A.

OpenStudy (rowanmartinfan):

i did...\[\sqrt{T^2} = \sqrt{a^3}\]

OpenStudy (rowanmartinfan):

then i came up with 3/2

OpenStudy (anonymous):

3/2 is not the answer, it is the exponent. Your square root operation was correct. But...\[\sqrt{T^{2}}=\sqrt{A^{3}}\]\[T=A^{3/2}\]

OpenStudy (rowanmartinfan):

ooh i did do it correct but i forgot the A = T ooh ok thanks

OpenStudy (anonymous):

And also \[A=T^{2/3}\] Now you need to solve for A in the equation with the K in it.

OpenStudy (anonymous):

I'm sorry, solve for T.

OpenStudy (rowanmartinfan):

i was about to say where did that K come from

OpenStudy (anonymous):

It does have a K in it, remember your problem? K is the multiple for the distance of planet Y. The equation is\[T^{2}=K^{3}A^{3}\]

OpenStudy (anonymous):

Solve for T, which is the orbital period

OpenStudy (rowanmartinfan):

oh i for got

OpenStudy (rowanmartinfan):

ooh i have no clue how to go about that with the ^3 thing

OpenStudy (anonymous):

\[T^{2}=K^{3}A^{3}\]\[\sqrt{T^{2}}=\sqrt{K^{3}A^{3}}=\sqrt{K^{3}}\sqrt{A^{3}}\]\[T=K^{3/2}A^{3/2}\]

OpenStudy (anonymous):

So for planet X, the orbital period, T, would be:\[T=A^{3/2}\] And for planet Y the orbital period would be:\[T=K^{3/2}A^{3/2}\]

OpenStudy (rowanmartinfan):

oh i see

OpenStudy (anonymous):

So by what factor is planet Y's orbital period larger than that of planet X?

OpenStudy (rowanmartinfan):

ok im back

OpenStudy (anonymous):

Okay. All you need to do is interpret the equations to get your answer.

OpenStudy (rowanmartinfan):

would u add both the tops & bottums like 3+3 & 2+2

OpenStudy (anonymous):

The question is basically asking you, "What would you multiply the orbital period of X by to get the orbital period of Y?"

OpenStudy (anonymous):

The orbital period of X is A^{3/2} and the period of Y is K^{3/2}A^{3/2}